An archer shoots 12 arrows at a target, with 9 arrows hitting the scoring area and 3 missing the scoring area. Write the ratio of the number of hits to the number of misses in three ways

Answer by stanbon(75887)

An archer shoots 12 arrows at a target, with 9 arrows hitting the scoring area and 3 missing the scoring area. Write the ratio of the number of hits to the number of misses in three ways
--------------------------------------------
9/3 = 3/1= 3:1
=====================
Cheers,
Stan H.